### v1

# Dense Video Object Captioning from Disjoint Supervision

Abstract
We propose a new task and model for dense video object captioning – detecting, tracking, and captioning trajectories of all objects in a video. This task unifies spatial and temporal understanding of the video, and requires fine-grained language description. Our model for dense video object captioning is trained end-to-end and consists of different modules for spatial localization, tracking, and captioning. As such, we can train our model with a mixture of disjoint tasks, and leverage diverse, large-scale datasets which supervise different parts of our model. This results in noteworthy zero-shot performance. Moreover, by finetuning a model from this initialization, we can further improve our performance, surpassing strong image-based baselines by a significant margin. Although we are not aware of other work performing this task, we are able to repurpose existing video grounding datasets for our task, namely VidSTG and VLN. We show our task is more general than grounding, and models trained on our task can directly be applied to grounding by finding the bounding box with the maximum likelihood of generating the query sentence. Our model outperforms dedicated, state-of-the-art models for spatial grounding on both VidSTG and VLN.
1 Introduction
Powered by gigantic datasets and models, language is becoming the output modality of the most capable artificial intelligence models brown2020language ; alayrac2022flamingo ; ouyang2022training ; li2023blip . Language unifies different tasks with the same output space raffel2020exploring ; chen2022pali , is more descriptive than discrete class labels wu2022grit ; long2023capdet , and naturally facilitates zero-shot prediction of novel tasks radford2021learning ; brown2020language . Inspired by advances in natural language understanding, the vision community has explored language in a number of tasks including image captioning chen2015microsoft , dense image captioning krishna2017visual , question answering antol2015vqa , video captioning Monfort_2021_CVPR and representation learning radford2021learning . However, likely due to the scarcity of large-scale, aligned training data, we are not aware of any existing single vision-language model that unifies both fine-grained spatial- (by detecting objects) and temporal- (by reasoning across time in videos) recognition.
In this paper, we propose a new task, model, and benchmark for dense video object captioning (Dense VOC) – the task of generating captions of object trajectories in a given video (Fig. 1). We design an end-to-end model that detects objects in a video, tracks them throughout the video, and also generates a textual description of each object trajectory. Specifically, our model first produces class-agnostic region proposals separately for each frame, followed by an association-based tracking module to group objects into trajectories. Features are then sampled from the object trajectory and fed into an autoregressive language decoder to generate captions for each trajectory (Fig. 2). Our task is therefore a superset of independent tasks commonly studied in vision – namely object detection everingham2015pascal ; lin2014microsoft , multi-object tracking dendorfer2021motchallenge ; dave2020tao and captioning chen2015microsoft ; Monfort_2021_CVPR . Whilst one could utilize separate models for each task, end-to-end trained models such as ours can produce more temporally coherent captions and capture object motions and object interactions. For example, our model predicts a global caption per trajectory, which significantly reduces “caption switches” compared to captioning per frame.
Figure 1: Overview of the dense video object captioning (Dense VOC) task. Given a video, we predict object trajectories (identities are denoted by colors) and their natural language description. Our task involves object detection, multi-object tracking, and captioning. We show our model’s output on an example from the VidSTG zhang2020does validation set.
However, datasets with captioned object trajectories are scarce. We show how we can train our model without any full annotations by using a mixture of disjoint tasks and datasets which supervise different parts of our model. In particular, we use COCO lin2014microsoft for object detection and tracking zhou2022global , Visual Genome krishna2017visual for dense image captioning, and Spoken Moments in Time Monfort_2021_CVPR for video captioning. We demonstrate a model trained on these disjoint datasets can perform our task zero-shot without access to any full, captioned object trajectories during training. Furthermore, it can serve as a powerful initialization for finetuning on our full task too.
Although we are the first to our knowledge to study the Dense VOC task, we can repurpose existing video grounding datasets for evaluation and domain-specific finetuning. VidSTG zhang2020does and Video Localized Narratives (VLN) voigtlaender2023connecting , originally designed for spatio-temporal sentence grounding – the task of finding the object tube in the video given a sentence query – contain (non-exhaustive) annotations for captioned object trajectories. We repurpose these annotations for our task by not using the sentence queries as inputs, but rather as expected model outputs. This means our task is more general and harder than the original grounding task, and we show that models trained on our task can be directly applied to grounding by selecting the bounding boxes with the maximum likelihood of generating the query sentence. We also develop evaluation metrics for our full task, which evaluate the captioning, detection and tracking abilities of our model.
Experiments show that our end-to-end trained Dense VOC model outperforms a recent, strong per-frame baseline wu2022grit followed by an offline tracker by a substantial margin, especially on the temporal consistency of the captions. Moreover, we achieve significant improvements from our disjoint, multi-dataset training. Furthermore, by directly applying our generative captioning model to the discriminative grounding task, we are able to outperform dedicated spatial grounding models on both VidSTG zhang2020does and VLN voigtlaender2023connecting datasets. We will release code at https://github.com/google-research/scenic.
2 Related Work
Image captioning chen2015microsoft ; anderson2018bottom ; xu2015show ; rennie2017self describes the content of an image with language. State-of-the-art methods map the input image to output text by using large end-to-end transformer models jiang2020defense ; desai2020virtex ; li2020oscar ; zhang2021vinvl ; li2023blip ; yu2022coca pretrained on large datasets sharma2018conceptual ; radford2021learning . For example, GIT wang2022git simply concatenates vision tokens from a ViT encoder dosovitskiy2020image with a begin-of-sentence token and generates texts using an auto-regressive language decoder vaswani2017attention ; devlin2018bert . Similar architectures apply to video captioning xu2016msr ; zhou2018towards ; Monfort_2021_CVPR , by concatenating wang2022git or pooling yan2022video features from each frame, before feeding them to the auto-regressive text decoder. Our work builds on existing captioning architectures wang2022git , and we train on video captioning datasets Monfort_2021_CVPR as a source of weak supervision for object trajectory captioning.
Dense object captioning in contrast aims to detect objects in an image and describe them with text johnson2016densecap ; li2019learning ; yin2019context ; shao2022region ; wu2022grit . It was popularized by the Visual Genome krishna2017visual dataset, which contains full annotations for the task. DenseCap johnson2016densecap uses a one-stage detector redmon2016you followed by an LSTM decoder hochreiter1997long on dense feature maps. Most recently, GRiT wu2022grit inherited the state-of-the-art image captioning architecture of GIT wang2022git , and generated object captions, also with a transformer decoder vaswani2017attention , from RoI-pooled he2017mask image features. Our model builds upon the approach of GRiT wu2022grit as a per-frame baseline, and extends it to videos via an end-to-end tracking and video-based training. We additionally note that dense video captioning in the literature refers to the task of localizing and captioning multiple temporal events in a video krishna2017dense ; zhou2018towards ; wang2021pdvc ; yang2023vid2seq . Our task, in contrast, requires tracking and captioning objects in a video, and therefore also requires spatial localization (similar to dense object captioning in images), which is why we name our task “dense video object captioning”.
Object tracking aims to detect objects and track them with a consistent identity label. The predominant approach is tracking-after-detection Bewley2016_sort ; zhang2021fairmot ; du20211st , i.e. first running detectors on each frame and then using a separate tracker to link them. While this works well for existing benchmarks with only a few classes dendorfer2021motchallenge ; Geiger2012CVPR ; Yang2019vis , it is more challenging in our case: we need tracks before captioning to have consistent outputs. Thus, our work follows end-to-end tracking cheng2021mask2former ; li2022videoknet ; wang2021end ; zhou2022global . We adopt a global tracker GTR zhou2022global , which handles tracking as pairwise association among all objects within a video clip. However, in contrast to GTR which focuses only on tracking, we also caption the trajectory.
Figure 2: Overview of our model. Our model first produces region proposals per-frame using a class-agnostic detector (left, trained with a detection loss $L_{object}$). We then run a global association module that group objects into trajectories (middle, trained with an association loss $L_{assoc}$). The trajectory features are then fed into a language decoder to produce a caption (right, trained with a caption loss $L_{caption}$). The language decoder can take both a single-frame feature or a trajectory feature as input. We (pre)train these modules on different and disjoint datasets (Sec. 3.4) and show this model can be applied zero-shot to our full task, or as initialization for finetuning.
Video object grounding zhang2020does ; voigtlaender2023connecting aims to find a spatio-temporal tube given a video and a query sentence input. Existing methods zhang2020does ; yang2022tubedetr ; jin2022embracing ; su2021stvgbert approach this task in a discriminative fashion: they first encode the query sentence and then use the sentence feature to find the corresponding object. Although our model is not trained for this task, we show it can still be used for grounding. As our model is generative, we use our language decoder to compute the likelihood of an object producing the query sentence, selecting the highest scoring one to perform grounding. Finally, we also note that these tasks are also related to video-referring object segmentation bellver2020refvos ; wu2022language ; yu2016modeling which grounds textual queries to segmentation masks of the input. Segmentation, however, is not the focus of this work.
3 Method
Our model builds upon prior techiniques for captioning and dense object captioning in images wu2022grit ; wang2022git which we first review in Sec. 3.1. As shown in Fig. 2, our model consists of region proposal, end-to-end tracking (Sec. 3.2), and captioning of the resulting trajectories (Sec. 3.3). We describe how we effectively train our model from disjoint tasks which supervise different parts of our model in Sec. 3.4. Finally, we explain how to apply our model to video object grounding in Sec. 3.5.
3.1 Background
The task of image captioning maps an input image, $\mathbf{I}^{H\times W\times 3}$, to a caption $c=(y_{1},y_{2},\ldots,y_{n_{t}})$ which is a sequence of up to $n_{t}$ text tokens from a given vocabulary. The minimal set of components is an image encoder, followed by a text decoder vaswani2017attention . The encoder maps the input image $\mathbf{I}$, to a feature representation $\mathbf{f}\in\mathbb{R}^{n_{v}\times d}$ consisting of $n_{v}$ tokens with dimensionality $d$. The subsequent text decoder is auto-regressive vaswani2017attention ; graves2013generating – it predicts the next text token, $y_{i}$, as a function of both the image features, $\mathbf{f}$, and previously generated text tokens, $\mathbf{y}_{1:i-1}$, denoted by $y_{i}=\text{Decode}(\mathbf{f},\mathbf{y}_{0:i-1})$. Note that the first step of decoding begins with $y_{0}=\texttt{BOS}$, a special beginning-of-sentence token, and the caption ends when the end-of-sentence token, EOS, is output by the model. This simple image captioning model has been demonstrated to be effective and scalable by GIT wang2022git , achieving state-of-the-art results across a number of captioning datasets.
GRiT wu2022grit extends the approach further to dense object captioning of images: Here, the authors use a region proposal network zhou2019objects to produce a set of $K$ class-agnostic bounding boxes, $b_{1},b_{2},\ldots,b_{K}$. Features corresponding to each of these regions are obtained using RoIAlign he2017mask , resulting in a localized feature, $f_{k}=\mathbb{R}^{r\times r\times d}$ where $r=7$ is the output resolution of RoIAlign. Each of these grid features is flattened into $f_{k}\in\mathbb{R}^{r^{2}\times d}$ and decoded independently by the text decoder, as done in GIT wang2022git . Therefore, the loss used to train a GRiT model consists of $L=L_{object}+L_{caption}$ where $L_{caption}$ is a cross-entropy loss over all text tokens in the vocabulary, and $L_{object}$ consists of bounding box regression and objectness terms, as standard in object detection literature zhou2019objects .
Next, we extend object captioning to videos, by tracking proposals over time, aggregating trajectory features, and captioning trajectories, in an end-to-end fashion.
3.2 End-to-end tracking
As shown in Fig. 2 (left), we first produce object proposals separately for each frame. Tracking aims to assign each object in each frame a unique trajectory identity $\delta\in\mathbb{N}$. We define $\mathbf{f}^{t}_{k}\in\mathbb{R}^{r^{2}\times d}$ as the ROI feature of object $k$ in frame $t$, $\mathbf{F}=\{\mathbf{f}^{t}_{k}\}_{t=1,k=1}^{T,K_{t}}$ as the concatenation of all object features in the video, and $M=|\mathbf{F}|=\sum_{t=1}^{T}K_{t}$, where $K_{t}$ is the number of objects at the $t^{\text{th}}$ frame.
From these object features $\mathbf{F}$, our network predicts a global association matrix, $\mathbf{A}\in\mathbb{R}^{M\times M}$, where $\mathbf{A}_{ij}=1$ if the objects denoted by the $i^{th}$ row and $j^{th}$ column respectively, are from the same trajectory at different time steps (as shown in Fig. 2 middle). Otherwise, $\mathbf{A}_{ij}=0$ meaning that these objects are from different trajectories, or at least one of them is background.
We use a transformer $\mathcal{F}$ with two self-attention layers, similar to GTR zhou2022global , to predict the association matrix $\mathbf{A}=\sigma(\mathcal{F}(\mathbf{F})$), where $\sigma$ is the sigmoid activation. For the ground truth supervision, we construct $\mathbf{\bar{A}}$ where $\mathbf{\bar{A}}_{ij}=1$ if and only if the row $i$ and column $j$ of $\mathbf{A}$ are matched to the same ground truth trajectory using an Intersection over Union (IoU) criteria of 0.5. The training loss $L_{assoc}$ for this module is then a binary cross entropy between $\mathbf{A}$ and $\mathbf{\bar{A}}$, $L_{assoc}=\frac{1}{M}\sum_{ij}\text{BCE}(A_{ij},\bar{A}_{ij})$.
To obtain the object identities $\{\delta_{k}^{t}\}_{t=1,k=1}^{T,K_{t}}$, needed for tracking evaluation, we design a greedy grouping algorithm taking $\mathbf{A}$ as input. Our algorithm greedily extracts the longest trajectories from untracked objects, until there are no possible associations (indicated by the association score being above a threshold $\theta$), and guarantees each trajectory has at most one object in each frame. We include the algorithm diagram in Appendix B.1, noting that it runs efficiently on GPU/ TPUs during training.
3.3 Trajectory captioning
Given the pairs of object features and identities ($\mathbf{f}_{k}$, $\delta_{k}$), we consider two strategies to caption them:
Hard aggregation. Let $\mathbf{f}_{\tau}=\{\mathbf{f}_{k}\}_{k\in\{\delta_{k}=\tau\}}$ be the set of all object features with identity $\tau$, we want to generate a single trajectory caption from $\mathbf{f}_{\tau}$. As a video can be long, it is expensive to naively use the entire $\mathbf{f}_{\tau}$. Inspired by the video captioning model of GIT wang2022git , we uniformly sample a subset of frames in the trajectory. Specifically, we denote this as $\mathbf{g}_{\tau}=\text{UniformSample}(\mathbf{f}_{\delta},m)$, where $\mathbf{g}_{\tau}\in\mathbb{R}^{m\times r^{2}\times d}$, where $m$ is the number of sampled frames. We set $m=6$ following ablations in Appendix C.1. We then autoregressively predict the output caption $\mathbf{y}$ as in Sec. 3.1, where $y_{i}=\text{Decode}(\mathbf{g},\mathbf{y}_{0:i-1})$. Note that the language decoder has the same parameters as single-frame object captioning, but processes more input tokens. As a result, it is trained in the same manner with a softmax cross-entropy loss over the vocabulary of text tokens, denoted by $L_{caption}$.
Soft aggregation. An alternative to concatenating features of a track as done above, is to compute a weighted sum of features in the same track from other frames. In fact, the association matrix $\mathbf{A}$ already serves as a trajectory feature aggregation weight. We set $\mathbf{F}^{\prime}=\frac{\mathbf{A}}{||\mathbf{A}||}\cdot\mathbf{F}$, where $\cdot$ denotes matrix multiplication, and |||||$\cdot$||| | ⋅ | | normalizes $\mathbf{A}$ by rows. Each row in $\mathbf{F}^{\prime}$, $\mathbf{f}_{k}$, is an object feature smoothed by its trajectory. We then feed each $\mathbf{f}_{k}$ to the language decoder and generate caption $\mathbf{y}$. This smoothes object features in a frame using other frames in the trajectory. Again, we do not introduce new parameters and use the same language decoder and loss $L_{caption}$.
3.4 Pretraining with disjoint subtasks
As shown in Fig. 2, our model is trained with the loss function, $L=L_{object}+L_{assoc}+L_{caption}$. Note that to avoid additional hyperparameters, we do not weight each loss term separately. Here we describe how we can decompose our full Dense VOC task into subtasks, and train each on different datasets, as shown in Tab. 1. This enables us to leverage more data, and to also perform our final task in a zero-shot manner (i.e. without any full dense video object captioning annotations).
Object detection. Using detection datasets for images, we can train the region proposal generator of our model by $L_{object}$. We use COCO lin2014microsoft as it is the most popular dataset for this task.
Dense captioning in images. Dense object captioning datasets of images allow us to train both the region proposal generator and the text decoder, by supervising $L_{object}$ and $L_{caption}$. Here, we use Visual Genome krishna2017visual , as it contains dense annotations for different regions. Note that these regions can include “stuff” forsyth1996finding classes like “sky”, or parts of objects, like clothes on a person. Although this enriches the object types we can detect, it also differs from the vocabulary of other datasets, introducing a domain shift.
Global video captioning. Video captioning datasets help us to reduce the domain gap to our final task by also training on video. In particular, we use Spoken Moments in Time (SMiT) Monfort_2021_CVPR which is the largest dataset for this task and contains narrations for short clips (roughly 3 seconds). As there are no object annotations, but only video-level captions, we construct a region proposal from the entire frame zhou2022detecting and caption that with our text decoder, applying the same $L_{caption}$. Note that our tracking module (Sec. 3.2) is effectively an identity function as we only have a single region proposal per frame. This approach is inspired by previous works in weakly-supervised detection which have trained object detectors from image-level or video-level supervision zhou2022detecting ; bilen2016weakly ; arnab2020uncertainty .
Tracking. Training the tracking module of our network (Sec. 3.2) requires annotations that associate detections of an object identity throughout the video. We found that existing tracking datasets either have too limited vocabularies for general objects (MOT dendorfer2021motchallenge , KITTI Geiger2012CVPR , YouTube VIS Yang2019vis ), or are too small (TAO dave2020tao and UVO wang2021unidentified label 600 and 5000 videos respectively). As a result, following existing work zhang2021fairmot ; zhou2022global ; zhou2020tracking , we instead augment image datasets into tracking ones by applying two different data augmentations to the same image, and then linearly interpolating the frames in between to form a pseudo-video. This allows us to apply $L_{assoc}$ and $L_{object}$ when training our model. In particular, we augment COCO (referred to as Aug-COCO), following CenterTrack zhou2020tracking .
Dataset Annotation type Train set size ($10^{3}$) $L_{object}$ $L_{assoc}$ $L_{caption}$
COCO lin2014microsoft Image detection 118
VG krishna2017visual Image object captioning 70
SMiT Monfort_2021_CVPR Video captioning 480
Aug-COCO lin2014microsoft Video object tracking 118
Table 1: Datasets for pretraining. We supervise different losses based on available annotations.
3.5 Application to video object grounding
The task of video object grounding zhang2020does ; voigtlaender2023connecting consists of two inputs: a video, $\mathbf{V}$, and a sentence query, $\bar{c}$. The output is a sequence of bounding boxes, $\{b^{s},b^{s+1},\dots,b^{e}\}$, corresponding to the sentence query, where $s$ and $e$ are the indices of the start and end frames respectively.
Our model, however, generates captions, $c$, at the output, rather than accepting it as an input. To apply our model to grounding, we follow an analogous approach to previous works which have performed closed-set image classification with captioning models alayrac2022flamingo ; chen2022pali : we evaluate the likelihood (i.e., exponential negative cross-entropy loss) of the sentence query, $\bar{c}$, for each of the object trajectories produced by our model. In practice, we find that instead of just taking the object trajectory with the highest sentence-likelihood, we achieve higher accuracy by weighting the likelihood by the detection score. Concretely, we run the proposal and tracking modules of our model to obtain object bounding boxes, detection scores, features, and track ids as $\{(b_{k}^{t},s_{k}^{t},\mathbf{f}_{k}^{t},\delta_{k}^{t})\}_{t=1,k=1}^{T,K_{t}}$. We choose the box with the highest weighted likelihood: ${k}^{*}=\text{argmax}_{k}(s_{k}^{t}\cdot\text{exp}(-L_{caption}(\text{Decode}(\mathbf{f}_{k}^{t}),\bar{c})))$, $b^{t}=b_{{k}^{*}}^{t}$.
4 Experimental Evaluation
4.1 Experimental setup
As we are proposing a new task, there is no dedicated dataset or evaluation metrics for dense video object captioning. Fortunately, existing video grounding zhang2020does ; voigtlaender2023connecting datasets have annotations for object trajectories and their captions, allowing us to repurpose them for Dense VOC, as defined below.
4.1.1 Datasets
VidSTG zhang2020does is a dataset for spatio-temporal sentence grounding, but can be used for our dense video object captioning task: Each video annotates multiple textual queries and their corresponding spatio-temporal tubes. By aggregating these across all videos, we obtain the paired trajectory-caption annotations that we need for training and evaluating our model. VidSTG has exhaustive trajectory (i.e. bounding box and tracking) annotations for all objects shang2019annotating , but not all objects are used in grounding, and thus not all objects have captions. We account for these partial annotations in both training and testing. Specifically, we do not compute $L_{caption}$ on objects without caption annotations, and also exclude them during evaluation (see Sec. 4.1.2). In particular, when a prediction is matched to a ground truth without caption annotations, we do not evaluate its captioning metrics, but still evaluate detection metrics. The dataset contains 5 436 training videos and 602 validation videos, with each video being at most 200 frames long. We use the declarative annotations from the dataset, containing 19K captioned trajectories (out of 28K total trajectories) and 2K captioned trajectories (out of 3K total trajectories) for training and testing respectively.
Video Localized Narratives (VLN) voigtlaender2023connecting is a recent dataset that asks annotators to narrate the “main actors” in a video. We therefore use these narrations as our target captions. Similar to VidSTG, these captions are not exhaustive for all objects, and so we exclude objects without captions in both training and evaluation of the captioning module. Importantly, VLN was created by adding narrations to existing video datasets. We use the VLN subset extending the UVO dataset wang2021unidentified which has exhaustive detection and tracking annotation for all objects wang2021unidentified . Each video in this subset has bounding box annotations for 3 sparsely sampled frames, and we thus train and evaluate on these annotated 3 frames. The dataset contains a total of 5 136 training and 2 451 validation videos, 9 917 training and 5 325 validation total trajectories, and 5 588 training and 3 071 validation captioned trajectories.
4.1.2 Evaluation metrics
[…]
Our task consists of captioning, detection and association. Therefore, we also define a “Captioning Accuracy” (CapA) term as $\text{CapA}=\frac{1}{3|TP^{\prime}|}\sum_{(x,y)\in TP^{\prime}}(\text{METEOR}(x,y)+\text{CIDEr}(x,y)+\text{SPICE}(x,y))$, where METEOR, CIDEr, and SPICE are three popular image-captioning metrics chen2015microsoft , and $TP^{\prime}$ are the true-positives detection pairs that have caption annotations (as discussed in Sec. 4.1.1). The final metric is then $\text{CHOTA}=\sqrt[3]{\text{DetA}\cdot\text{AssA}\cdot\text{CapA}}$, effectively adding a captioning term to the HOTA metric. We include further implementation details and evaluation code in the supplementary.
Frame mAP-METEOR ($\text{AP}_{M}$). We additionally reuse the evaluation metric for dense object captioning in images from Visual Genome krishna2017visual . Such frame-based metrics evaluate the model’s predictions given the temporal context of the video, and are also popular in spatio-temporal action detection gu2018ava ; soomro2012ucf101 . Note that we avoid the captioning threshold for objects without captioning annotations, as detailed in Appendix B.2.
[…]
4.1.3 Implementation details
We implement our model in JAX jax2018github using the scenic libinary dehghani2021scenic . We use GRiT wu2022grit as our baseline. GRiT uses a ViTDet-Base dosovitskiy2020image ; li2022exploring backbone (initialized with CLIP radford2021learning ), a CenterNet zhou2019objects proposal network and RoI Head, and a randomly-initialized text decoder following that of GIT wang2022git ; vaswani2017attention .
We first train our model for general Dense VOC on large-scale disjoint datasets (Sec. 3.4). During disjoint pretraining, we sample batches from different datasets with an even ratio, (1: 1: 1: 1), thus avoiding additional hyperparameters. For video datasets, we sample 8 frames for each video and use a local batch size of 1. For image datasets, we use a local batch size of 8. We train our model on 32 TPU cores, which means we have an effective batch size of 256 images or 32 videos.
We then evaluate on our two fully-annotated datasets (Sec. 4.1.1) in both zero-shot and full-finetuning setups. For VidSTG, we sample 16 frames during training, and then run on all 200 frames during testing. For VLN, we use all 3 annotated frames in both training and evaluation. In all cases, we use a frame resolution of $384\!\times\!384$. We include exhaustive implementation details and hyperparameters in Appendix B.4 and commit to releasing code and models.
4.2 Main Results
# COCO VG SMiT Aug- VidSTG VLN
COCO CHOTA DetA AssA CapA $\text{AP}_{M}$ CHOTA DetA AssA CapA $\text{AP}_{M}$
1 - 50.5 - - - - 27.9 - - -
2 - 18.9 - 9.4 18.7 - 12.2 - 7.2 10.1
3 - - - - - - - - - -
4 - 20.8 - 10.2 19.2 - 14.1 - 9.1 12.2
5 - 41.5 - 9.4 38.3 - 26.7 - 7.7 19.4
6 - 51.7 - 4.0 35.2 - 28.7 - 6.4 15.9
7 - 45.1 - 9.8 39.0 - 24.4 - 8.2 19.9
8 28.0 45.9 48.0 10.0 39.8 28.2 28.6 83.4 9.4 21.3
Table 2: Zero-shot evaluation of our disjoint trained models with varying datasets. We show results on the VidSTG zhang2020does (center block) and VLN voigtlaender2023connecting (right block) datasets without finetuning on them. For models without tracking supervision (#1–7), we cannot report their association accuracy (AssA). Our full model (#8) gains full Dense VOC ability from disjoint training, and shows good performance on all metrics. More detailed captioning metrics are in Appendix C.3.
Zero-shot evaluation. We first pretrain on multiple disjoint datasets (Sec. 3.4), and evaluate zero-shot on our target datasets, VidSTG and VLN, without training on them in Tab. 2. Note that zero-shot evaluation is trivial to perform for models that directly generate text compared to classification, thanks to their open-vocabulary nature. There are, however, significant domain gaps in the object classes, captioning style, and video durations between our training and evaluation datasets.
As mentioned in Tab. 1 and Fig. 2, each dataset supervises different parts of our model. For example, a model that is only trained on COCO (#1), is only trained with $L_{object}$, meaning that it only produces region proposals which we can evaluate with the Detection Accuracy component of our CHOTA metric. Similarly, Visual Genome (VG) supervises both the region proposal and captioning heads of our model. However, there is a large domain gap between the captions in VG, and our target datasets, since the captions in VG are for single images and focus on very different vocabularies. Moreover, as VG tends to annotate bounding boxes around object parts rather than entire objects, the Detection Accuracy in is lower than a model trained on COCO alone (comparing #2 to #1). The domain gap remains with SMiT, as although it captions videos, they are for the entire clip and not at an object level. Note that we cannot evaluate a model trained only on SMiT: it does not produce bounding boxes as there is no supervision for it.
We observe in Tab. 2 that the different datasets have complementary properties: Adding COCO improves detection accuracy (#2 to #5, #4 to #7), and adding SMiT improves the captioning accuracy (#2 to #4, #5 to #7) even though SMiT only captions at a global, video-level. Finally, training with Aug-COCO allows us to also supervise $L_{assoc}$ and thus the tracking module of our model. A model trained on all of our pretraining datasets (#8) can therefore perform the full Dense VOC task, and shows good performance on all individual metrics compared to models trained on fewer datasets.
# COCO VG SMiT Aug- VidSTG VLN
COCO CHOTA DetA AssA CapA $\text{AP}_{M}$ CHOTA DetA AssA CapA $\text{AP}_{M}$
0 49.3 61.2 54.8 35.8 64.1 29.5 35.4 82.5 8.8 35.1
1 51.6 65.6 56.9 36.8 69.3 31.4 43.9 86.2 8.2 36.3
2 52.3 64.6 58.4 38.0 68.7 39.8 45.1 84.2 16.6 45.9
3 46.9 60.3 40.0 33.0 59.3 36.1 41.2 78.9 14.5 38.0
4 52.7 64.4 58.1 39.1 70.0 40.2 44.2 83.4 17.6 47.2
5 54.3 66.6 61.0 39.6 71.2 40.4 44.1 86.6 17.2 46.1
6 51.9 65.2 57.7 37.1 69.6 35.3 43.7 86.1 11.6 41.3
7 54.5 66.1 60.2 40.7 71.2 40.8 44.2 87.0 17.6 48.2
8 54.3 66.1 59.8 40.5 71.2 40.8 44.3 86.9 17.6 48.2
Table 3: Finetuning evaluation of our disjoint trained models with varying pretraining datasets.. We show results on VidSTG zhang2020does (center block) and VLN voigtlaender2023connecting (right block). Each row is a model pretrained on the specified datasets and then finetuned on the downstream datasets. #0 is finetuned from a CLIP radford2021learning checkpoint. Our disjoint-pretrained models perform better than single-dataset pretrained ones, and yield the best CHOTA and $\text{AP}_{M}$ on both VidSTG and VLN. The best entries are bolded, and the second-best underlined. More detailed caption metrics are in Appendix C.3.
Finetuning evaluation. We now finetune each of the pretrained models from Tab. 2 and show results in Tab. 3. We also include a baseline (#0) which initializes from only a CLIP-pretrained checkpoint radford2021learning , observing that this model performs the worst. If we pretrain on only a single dataset, we find that Visual Genome (VG) is the best, probably because it allows us to pretrain both the captioning- and region proposal modules (Fig. 2). Once again, we observe that different pretraining datasets are complementary, as adding either SMiT or COCO (#2 to #4, and #2 to #5) improves results further. We achieve the best results with our model pretrained on COCO, VG and SMiT (#7), which outperforms the best single-dataset pretrained model by 2.2 CHOTA on VidSTG, and 1 CHOTA on VLN. The improvement over only a CLIP-pretrained checkpoint is even larger, by 5.2 CHOTA and 11.3 CHOTA on the two respective datasets. We did not observe any improvement from including Aug-COCO during pretraining when finetuning on VidSTG or VLN, although this is important for zero-shot performance. This is likely because the model can learn tracking sufficiently from the finetuning annotations alone. We provide qualitative visualizations in Fig. 3 and the supplement.
VidSTG VLN
CHOTA($\uparrow$) CapA($\uparrow$) CapSwitch($\downarrow$) CHOTA($\uparrow$) CapA($\uparrow$) CapSwitch($\downarrow$)
Per-frame with tracker 48.3 37.1 38.1% 39.2 15.6 33.4%
Our tracking (hard) 52.1 37.5 14.7% 39.8 16.6 7.6%
Our tracking (soft) 52.3 38.0 17.5% 39.5 16.2 15.7%
Table 4: Ablation of our tracking module. We show results on VidSTG zhang2020does (center block) and VLN voigtlaender2023connecting (right block). The models are based on #2 of Tab. 3 on the corresponding datasets. The baseline model in the first row is trained on the same datasets and runs an IoU-based tracker as a post-processing step. Our end-to-end tracking-and-captioning approach generates temporally smooth captions (better CapSwitch), and mildly improves captioning quality (better CapA).
Ablation of our tracking module. We next ablate the importance of our end-to-end tracking module. As a baseline, we use a per-frame model without our tracking module. Specifically, we pretrain it on Visual Genome, and then finetune it on VidSTG and VLN respectively whilst only processing a single frame. The tracking is done as a post-processing step, using an IoU-based tracker from Detectron2 wu2019detectron2 . As shown in Tab. 4, both our tracking approaches substantially improve the baseline on caption switches. On VLN, tracking is easier as there are only 3 frames, and all trackers perform similarly in CHOTA. We also observe that our hard- and soft-aggregation (Sec. 3.3) both mildly improve captioning quality, and that caption switches are improved further by hard aggregation. Based on these results, we used soft aggregation on VidSTG and hard aggregation on VLN.
4.3 State-of-the-art comparison on video grounding
As introduced in Sec. 3.5, models trained for Dense VOC can be directly used for sentence grounding, by finding the proposals with the maximum likelihood of generating the query sentence. We evaluate on the Location-QA and spatial grounding benchmarks of VLN voigtlaender2023connecting and VidSTG zhang2020does respectively.
Figure 3: Qualitative results on VidSTG. Our model successfully handles fast motion (1st row) and crowded scenes (2nd row). However it may misrecognize objects (2nd row, “dog” should be “goat”), and be inaccurate at action boundaries (3rd row, the caption is “holds hand” before it happens).
Table 5: State-of-the-art comparison of Location-QA on VLN voigtlaender2023connecting . We show the official metrics consisting of recall $>0.5$, precision $>0.5$, and the combination of the two. We compare to the baseline provided in VLN paper voigtlaender2023connecting , GRiT wu2022grit , and our full model (#8 of Tab. 2).
Recall Precision Both
ReferFormer wu2022language ; voigtlaender2023connecting 66.7 53.9 48.3
GRiT wu2022grit 77.4 68.5 62.1
Ours 86.7 68.4 65.1
Finetuned Zero-shot
STVGBert su2021stvgbert 47.3 -
TubeDETR yang2022tubedetr 59.0 -
STCAT jin2022embracing 61.7 -
Ours 61.9 54.1
Table 5: State-of-the-art comparison of Location-QA on VLN voigtlaender2023connecting . We show the official metrics consisting of recall $>0.5$, precision $>0.5$, and the combination of the two. We compare to the baseline provided in VLN paper voigtlaender2023connecting , GRiT wu2022grit , and our full model (#8 of Tab. 2).
Table 6: State-of-the-art comparison of spatial grounding on the VidSTG test set. All models assume ground truth temporal localization is provided.
VLN Location-QA consists of questions starting with “Where is”, and requires the model to produce a bounding box at each frame in the video. The task is therefore effectively a sentence grounding problem, and indeed, the ReferFormer wu2022language baseline used by voigtlaender2023connecting removes the “Where is” in the question, and performs sentence grounding. We adopt this approach too, as detailed in Sec. 3.5.
In this dataset, only one annotated frame (unknown at inference time) is evaluated, and this benchmark therefore effectively does not involve temporal localization. As the annotation of this dataset is based on mouse traces instead of bounding boxes, the evaluation metric considers bounding box coverage (recall), precision, and both in conjunction. We refer to the VLN paper voigtlaender2023connecting for full details on evaluation metrics. As shown in Tab. 6, we improve substantially over ReferFormer and our GRiT wu2022grit baseline. Note that none of the models are trained on the VLN Location-QA training split.
VidSTG requires producing a sequence of bounding boxes for a given sentence query. The evaluation metric is the average of the Intersection over Union (IoU) at each frame, between the predicted and ground truth bounding boxes for the target object. We compare to other prior works on this dataset in Tab. 6, assuming that the input video is already trimmed temporally to the objects of interest. Our model achieves the best IoU, outperforming models designed specifically for grounding, thereby showing that our generative framework can be used effectively in the discriminative grounding task. In contrast to prior work, we also evaluate zero-shot without training on VidSTG, and still perform competitively in this scenario. This result emphasizes the efficacy of our disjoint pretraining. We provide results with temporally untrimmed videos in Appendix C.4.
5 Conclusion
We proposed the new task of dense video object captioning. Although this task requires expensive annotations across space, time and language, we show we can train a model on existing larger-scale datasets on disjoint subtasks. Moreover, our proposed end-to-end architecture is important for producing more accurate and coherent captions. We discuss broader impacts in Appendix A.
Limitations. Our work is a first step towards fine-grained, object-level video and language understanding. Currently, our end-to-end model produces object trajectories first, and then generates a single caption for the entire detected trajectory (Sec. 3.3). Therefore, it cannot caption multiple action segments within a single object trajectory, and we aim to handle this in future work, building upon approaches for dense video captioning yang2023vid2seq . Similar to other video models, we cannot handle very long videos arnab2021vivit ; wang2021end ; wu2022language , and we only process up to 200 frames in this paper. Nevertheless, we hope our work opens an exciting avenue for future research in fine-grained video-language understanding. Acknowledgments. We thank Shyamal Buch and Shen Yan for helpful discussion and feedback.
Appendix A Broader impact
Our work presents a new task and model for dense video object captioning. This task represents a general technology with a wide range of potential applications. Whilst we are unaware of all potential applications of such models, it is important to be cognizant that each application has its own merits and societal implications depending on the intentions of the individuals building and using the system. For example, we believe that the Dense VOC models can be used as part of systems to improve video search and retrieval, though they could also be used in video surveillane systems too. Additionally, we note that training datasets, especially for captioning hendricks2018women ; zhao2021understanding , can contain biases that may render models trained on them unsuitable for deployment.
Appendix B Additional Experimental and Implementation Details
B.1 Tracking algorithm details
We use a greedy algorithm to decode an association score matrix into identities of objects (Sec. 3.2). To make the input shape consistent across training iterations, we zero-pad or truncate the number of objects in each frame $K_{t}$ to the same number $K=16$. We provide our trajectory ID assignment algorithm in Algorithm 1 below. The algorithm is implemented in JAX jax2018github , where the while-loop can run efficiently on GPU or TPU accelerators. We use $\theta=0.7$ throughout our experiments, where $\theta$ is the association score threshold for determining if two objects should be merged into the same track.
Input :  Number of frames T
 Number of objects per frame K
 Association Matrix $\bf{A}\in\mathbb{R}^{TK\times TK}$
Hyper parameters : Association score threshold $\theta$
Output :  Identities for each object $\delta\in\mathbb{N}^{TK}$
$M\leftarrow T\times K$ // Number of total objects.
$A\leftarrow$ preprocess($A$) // Preprocess $A$ to ensure objects in the same frame have a score of 00.
[…]
id_count $\leftarrow 0$ // Initialize ID count.
while $\hat{A}$.any() do
       track_len $\leftarrow$ $\hat{A}$.sum(axis=1) // Compute number of objects in each merge.
       i $\leftarrow$ track_len.argmax() // Find the longest track to merge.
       id_count $\leftarrow$ id_count + 1 // Create a new identity.
       $\delta\leftarrow\delta+$ id_count * $\hat{A}_{i}$ // Assign the current track a new ID using $\hat{A}_{i}$ as the binary mask.
       $\hat{A}\leftarrow\hat{A}-\hat{A}_{{i\cdot}}|\hat{A}_{{\cdot i}}$ // Remove merged indices from the binary matrix. “||||” is the logical-or.
end while
return $\delta$
Algorithm 1 Greedy assignment of identities from an association matrix.
B.2 $\text{AP}_{M}$implementation details
mAP-METEOR is the official evaluation metric used in Visual Genome krishna2017visual dataset for dense image object captioning. This metric evaluates predictions in each frame separately, without evaluating the tracking output.
mAP-METEOR is based on the Average Precision used in object detection lin2014microsoft ; everingham2015pascal , but includes a caption similarity criteria for determining true positives: i.e. a prediction is a true positive if the Intersection over Union (IoU) with the ground truth bounding box is above a threshold, and if the METEOR score banerjee2005meteor is above another threshold. We follow the same implementation and thresholds as the Visual Genome dataset [^1] https://github.com/jcjohnson/densecap/blob/master/eval/eval_utils.lua. i.e. IoU thresholds of (0.3, 0.4, 0.5, 0.6, 0.7) and METEOR thresholds of (0.0, 0.05, 0.1, 0.15, 0.2).
In our case, some objects in the datasets zhang2020does ; voigtlaender2023connecting only have bounding box annotations and no caption annotations (Sec. 4.1.1). For these objects without caption annotations, we allow any caption prediction (and therefore ignoring it) by setting it’s METEOR score to the maximum of 1. For brevity, we abbreviated this metric as $\text{AP}_{M}$in the main paper.
B.3 CHOTA Evaluation code
We will release our CHOTA evaluation code. This evaluation code is based on the official HOTA implementation [^2] https://github.com/JonathonLuiten/TrackEval. The original code is under an MIT license.
B.4 Full training details
As mentioned in Sec. 4.1.3, our model is based on GRiT wu2022grit . The original GRiT code [^3] https://github.com/JialianW/GRiT is released under an MIT license. Following GRiT, we use a ViTDet-Base dosovitskiy2020image ; li2022exploring backbone, a CenterNet zhou2019objects region proposal network and RoI Head, and a randomly-initialized text decoder following that of GIT wang2022git . The text decoder consists of 6 self-attention layers with casual feature masks wang2022git . All model architecture parameters follow the defaults from GRiT wu2022grit .
The original GRiT uses an MAE pretrained checkpoint, while in our case we found a CLIP pretrained checkpoint radford2021learning performs better on our task. To fit more frames into memory for both training and evaluation, we use a $384\!\times\!384$ input size instead of the original $1024\!\times\!1024$. This choice moderately decreases dense image captioning performance on Visual Genome (from 17.3 $\text{AP}_{M}$to 15.7 $\text{AP}_{M}$).
During disjoint multi-dataset pretraining, we sample batches from different datasets in an even ratio ($1:1:1:1$). For images datasets, a batch is composed of different images; for video datasets, we put the time dimension in batches and always guarantee images in the same mini-batch are from the same video. We use a local batch size of either 1 video (consisting of 8 sampled frames), or 8 images. As we use 32 TPU cores, this means that our global batch size is either 32 videos or 256 images. We use the AdamW optimizer with a learning rate of $2\times 10^{-4}$, weight decay of $0.05$, and a layerwise learning rate decay of $0.7$ li2022exploring ; wu2022grit . We train for $22.5\times 10^{3}$ iterations per dataset, decreasing the learning rate by a factor of 10 after $90\%$ and $97.5\%$ of the training schedule wu2022grit . For pretraining on all the 4 datasets in Sec. 3.4, this corresponds to a total of $90\times 10^{3}$ iterations, which took approximately 20 hours on 32, 16GB TPU cores.
For VidSTG zhang2020does finetuning, we sample 16 frames in training, and run on all 200 frames in testing. For VLN voigtlaender2023connecting finetuning, we use the 3 annotated frames in both training and evaluation. For finetuning experiments on both datasets, we use a video batch size $16$ and train for $11.25\times 10^{3}$ iterations, with a learning rate of $10^{-5}$, weight decay of $0.05$, and layerwise-learning decay of $0.7$ li2022exploring . The finetuning took approximately 6 hours on 16, 16GB TPU cores for VidSTG, and about 2 hours on 16, 16GB TPU cores for VLN. Inference on VidSTG requires 32GB of TPU memory to fit 200 frames.
Training losses. Training our model involves a detection loss $L_{object}$, a tracking loss $L_{assoc}$, and a captioning loss $L_{caption}$, that is
$L=L_{object}+L_{assoc}+L_{caption}.$ (1)
For completeness, we detail these three terms next:
The detection loss zhou2019objects involves a center heatmap loss, a bounding box regression loss, and a classification and bounding box refinement loss in the RoI head:
$L_{object}=L_{heatmap}+L_{reg}+L_{roi\text{-}cls}+L_{roi\text{-}reg}.$ (2)
The heatmap loss is defined on the predicted heatmap $Y\in\mathbb{R}^{H\!\times\!W}$ and the ground truth heatmap $\bar{Y}\in\mathbb{R}^{H\!\times\!W}$:
$L_{heatmap}(Y,\bar{Y})=\frac{1}{n}\sum_{ij}\begin{cases}(1-{Y}_{ij})^{\alpha}\log({Y}_{ij})&\!\text{if}\ \bar{Y}_{ij}=1\\ (1-\bar{Y}_{ij})^{\beta}({Y}_{ij})^{\alpha}\log(1-{Y}_{ij})&\!\text{otherwise,}\end{cases}$ (3)
where $n$ is the number of objects in the image, $\alpha=2$ and $\beta=4$ are the focal loss weights lin2017focal .
$L_{reg}$ is a gIoU loss union2019metric :
$L_{reg}(B,\bar{B})=\frac{1}{n}\sum_{i}(\text{IoU}(B_{i},\bar{B}_{i})-\frac{|C_{i}\backslash(B_{i}\cup\bar{B}_{i})|}{|C_{i}|}),$ (4)
where $B$ and $\bar{B}$ are the predicted and the ground truth bounding boxes of the $n$ annotated objects, $C_{i}$ is the enclosing convex hull of $B_{i}$ and $\bar{B}_{i}$, and ||$\cdot$|| ⋅ | computes the area.
$L_{roi\text{-}cls}$ is a softmax classification loss on each RoI box, defined on the predicted class logits $\textbf{c}\in\mathbb{R}^{2}$ and the ground truth label $\bar{c}\in\{0,1\}$. Here we only have foreground or background classification.
$L_{roi\text{-}cls}(\textbf{c},\bar{c})=-\log\text{softmax}(\textbf{c})_{\bar{c}}$ (5)
$L_{roi\text{-}reg}$ is an L1 loss between the predicted boxes $B$ and the ground truth boxes $\bar{B}$,
$L_{roi\text{-}reg}(B,\bar{B})=|B-\bar{B}|.$ (6)
The tracking loss is a per-element binary cross-entropy loss between the predicted association matrix $A$ and the ground truth binary matrix $\bar{A}$:
$L_{assoc}=-\frac{1}{M}\sum_{ij}(\hat{A}_{ij}\log{A_{ij}}+(1-\hat{A}_{ij})\log{(1-A_{ij})})).$ (7)
The captioning loss is a softmax on each predicted word over the entire vocabulary, with a label smoothing co-efficient of $0.1$ following GIT wang2022git .
$L_{caption}=\frac{1}{L}\sum_{i=1}^{L}\text{Smooth-Softmax-CE}(\text{Decode}(f,\bar{y}_{1:i-1}),\bar{y}_{i}),$ (8)
where $\bar{y}$ is the ground truth caption, $L$ is the groud-truth sentence length, and $f$ is the object feature.
Appendix C Additional Experimental Analysis
C.1 Ablation of hard tracking sampling.
We analyze the effect of the number of sampled frames, $m$, in hard-aggregation (Sec. 3.3) in Tab. 7. With hard-aggregation, the captioning accuracy benefits from a larger number of frames $m$, thanks to longer input-sequence length. However this also costs more TPU memory in both training and testing. We use $m=6$ in our ablation experiments (Tab. 4) as it achieves a good balance between accuracy and computation. Moreover, it also follows the default number of frames used in the GIT wang2022git video captioning model. On the other hand, soft-aggregation (i.e., weighted-sum features) does not have this hyperparameter and has a constant memory cost, and achieves a higher caption accuracy. However, it also results in a higher number of caption switches.
Num sampled frames ($m$) CHOTA($\uparrow$) DetA($\uparrow$) AssA($\uparrow$) CapA($\uparrow$) $\text{AP}_{M}$($\uparrow$) CapSwitch($\downarrow$)
1 51.4 66.1 59.0 34.8 68.9 15.8
2 53.0 66.1 59.2 38.1 70.6 15.9
4 53.3 66.2 59.5 38.4 70.9 14.7
6 53.5 66.2 59.4 39.0 71.2 14.5
8 53.7 66.2 59.4 39.4 71.2 12.9
Soft aggregation 54.3 66.1 59.8 40.5 71.2 17.5
Table 7: Hyper-parameter sweep for number of sampled frames, $m$, for hard-tracking. We show results on VidSTG zhang2020does validation set. The models are based on #8 of Tab. 3 on VidSTG. Results with hard-feature aggregation get improved with more sampled frames, given the cost of increased training and testing memory.
C.2 Using the UVO dataset for disjoint pretraining
For the disjoint pretraining of our model (Sec. 3.4), we used Augmented COCO as our tracking dataset. Another alternative would have been to use UVO wang2021unidentified , which contains real-world videos, but is relatively small at only 5000 videos.
Table 8 compares Aug-COCO and UVO under the setting of Tab. 2 and Tab. 3 #8, both using a default multi-dataset sampling ratio $1:1:1:1$. We observe that disjoint pretraining with Aug-COCO consistently performs better than UVO in both zero-shot and finetuning scenarios, thus motivating our choice to use Aug-COCO for our experiments in the main paper.
Tracking dataset CHOTA DetA AssA CapA $\text{AP}_{M}$
Zero-shot
UVO 24.8 41.2 52.1 7.1 33.8
Aug-COCO 28.0 45.9 48.0 10.0 39.8
Finetuning
UVO 50.9 65.2 53.4 37.9 70.4
Aug-COCO 54.3 66.1 59.8 40.5 71.2
Table 8: Results using UVO wang2021unidentified as the tracking dataset. We show both zero-shot results (top) and finetuning results (bottom) on VidSTG datasets. For reference, we also include our results of using Aug-COCO (#8 of Tab. 2 and Tab. 3). Aug-COCO performs better in both settings, motivating our choice to use it in the main paper.
C.3 Detailed captioning results
The Captioning Accuracy (CapA) component of our CHOTA metric is the average of the CIDEr, METEOR and SPICE metrics. For completeness, we report each of these captioning metrics individually in Tabs. 9 and 10, for zero-shot and full-finetuning evaluation, respectively.
# COCO VG SMiT Aug- VidSTG VLN
COCO CapA CIDEr METEOR SPICE CapA CIDEr METEOR SPICE
1 - - - - - - - -
2 9.4 7.1 8.1 13.1 7.2 2.6 7.2 11.7
3 - - - - - - - -
4 10.2 7.3 8.7 14.7 9.1 3.7 9.1 14.3
5 9.4 7.3 8.1 12.9 7.7 3.0 7.7 12.5
6 4.0 1.9 6.1 4.0 6.4 2.5 8.2 8.4
7 9.8 7.4 8.4 13.7 8.2 3.0 8.3 13.3
8 10.0 7.1 8.5 14.4 9.4 4.3 9.4 14.5
Table 9: Detailed captioning metrics of our zero-shot evaluation (Tab. 2). We show the individual captioning metrics CIDEr, METEOR, and SPICE for each row on both datasets.
# COCO VG SMiT Aug- VidSTG VLN
COCO CapA CIDEr METEOR SPICE CapA CIDEr METEOR SPICE
0 35.8 43.2 29.0 35.0 8.8 4.2 14.8 7.4
1 36.8 46.1 28.5 35.8 8.2 6.7 13.9 3.9
2 38.0 47.5 29.8 36.7 16.6 13.6 20.1 16.1
3 33.0 35.8 28.7 35.8 14.5 9.4 19.6 14.5
4 39.1 49.3 30.3 37.6 17.6 13.9 21.2 17.7
5 39.6 50.2 30.5 38.1 17.2 14.8 20.4 16.4
6 37.1 46.5 28.9 35.7 11.6 8.1 17.2 9.5
7 40.7 52.9 30.8 38.4 17.6 14.2 21.5 17.1
8 40.5 52.6 30.4 38.5 17.6 14.1 21.4 17.4
Table 10: Detailed captioning metrics of our finetuning evaluation (Tab. 3). We show the individual captioning metrics CIDEr, METEOR, and SPICE for each row on both datasets.
Validation set Test set
sIoU tIoU vIoU sIoU tIoU vIoU
STGRN zhang2020does - - - 38.0 48.5 19.8
STVGBert su2021stvgbert - - - 47.3 - 24.0
TubeDETR yang2022tubedetr 56.4 47.2 28.7 59.0 48.1 30.4
STCAT jin2022embracing - - - 61.7 50.8 33.1
Ours (zero-shot) 51.8 40.0 22.0 54.1 40.2 22.5
Ours 58.7 41.8 25.9 61.9 41.1 26.3
Table 11: State-of-the-art comparison of spatial-temporal grounding on VidSTG. “-” means the numbers are not reported in the paper. Our model performs competitively at this task, although it was not actually designed for it. As our model generates object trajectories without conditioning on the input query, it struggles at temporal localization, denoted by the tIoU. The spatial localization performance, denoted by the sIoU, outperforms dedicated methods for this task.
C.4 VidSTG spatio-temporal grounding evaluation
Table 6 of the main paper comapred to prior methods on the spatial-video grounding task on VidSTG (where the input videos were assumed to already be temporally trimmed). In Tab. 11, we report results for spatial-, temporal- and spatio-temporal grounding by reporting the Spatial IoU (sIoU), Temporal IoU (tIoU) and Video IoU (vIoU) respectively.
The sIoU assumes the video is already temporally trimmed before evaluation, thus evaluating spatial localization. Similarly, the tIoU assumes that the video is already cropped spatially around the object of interest, and only the temporal extent of the query sentence needs to be determined, thereby evaluating temporal localization. The vIoU that evaluates both spatial and temporal localization.
Our model was designed for the Dense VOC task, and not grounding, and we were able to perform grounding by selecting the bounding boxes with the highest likelihood of generating the target sentence (Sec. 3.5). As shown in Tab. 11, this approach works well for spatial-grounding, outperforming prior works in terms of the sIoU. However, as our model first generates object trajectories, without taking the input query into account, it struggles more at temporal localization. Nevertheless, it still achieves competitive results compared to prior works for both the tIoU and vIoU, although our model was not designed specifically for this task like the other methods in Tab. 11 which include explicit temporal localization modules within the network architecture.